# ChroniclingAmericaQA: A Large-scale Question Answering Dataset based on Historical American Newspaper Pages

### v1

@@ Abstract.
Question answering (QA) and Machine Reading Comprehension (MRC) tasks have significantly advanced in recent years due to the rapid development of deep learning techniques and, more recently, large language models. At the same time, many benchmark datasets have become available for QA and MRC tasks. However, most existing large-scale benchmark datasets have been created predominantly using synchronous document collections like Wikipedia or the Web. Archival document collections, such as historical newspapers, contain valuable information from the past that is still not widely used to train large language models. To further contribute to advancing QA and MRC tasks and to overcome the limitation of previous datasets, we introduce ChroniclingAmericaQA, a large-scale dataset with 485K question-answer pairs created based on the historical newspaper collection Chronicling America. Our dataset is constructed from a subset of the Chronicling America newspaper collection spanning 120 years. One of the significant challenges for utilizing digitized historical newspaper collections is the low quality of OCR text. Therefore, to enable realistic testing of QA models, our dataset can be used in three different ways: answering questions from raw and noisy content, answering questions from cleaner, corrected version of the content, as well as answering questions from scanned images of newspaper pages. This and the fact that ChroniclingAmericaQA spans the longest time period among available QA datasets make it quite a unique and useful resource.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: International Conference on Information Retrieval (SIGIR); July 14–18, 2024; Washington, USAbooktitle: SIGIR 2024isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Question answeringccs: Information systems Content analysis and feature selection
@@ 1. Introduction
However, leveraging these documents for QA is complex and poses a few challenges. The language used in heritage documents often differs significantly from the contemporary language in terms of vocabulary, syntax, and the overall context of the distant times (tah, 2021; Tahmasebi et al., 2021). These raises doubts if QA models trained on current text data can indeed perform well on documents from distant past. Additionally, the documents are often available only in scanned form, and the quality of the text obtained through Optical Character Recognition (OCR) can greatly vary, especially over long time periods, adding another layer of complexity to the task. Finally, these documents have a complex layout, often containing different font sizes and type throughout their content, which makes information extraction a complex task. Figure 1 shows examples of scanned newspaper pages and the OCRed text.
@@
Figure 1. An example of the scanned newspaper page from Chronicling America Collection. a) depicts the entire newspaper page published in Evening Star on 1803-02-07 in the District of Columbia, b) depicts the zoomed-in paragraph of the newspaper page shown in (a), c) shows the original OCR text of the zoomed-in paragraph that is available in the Chronicling America, and d) shows the OCR text corrected by GPT 3.5 Turbo.
@@
Despite the inherent challenges of using historical newspapers, they present an attractive and unique opportunity for research in the field of QA. The wealth of information in these documents makes them valuable resources that are not readily available in typical evaluation benchmarks. To address the challenges and utilize the potential of historical document collections, we introduce the ChroniclingAmericaQA, a large-scale dataset based on Historical American Newspaper Pages. As an underlying historical document collection, we use Chronicling America [^1] https://chroniclingamerica.loc.gov/about/, a digitized collection of America’s Historical newspaper pages from 1756-1963. Chronicling America is a project that was created with the aim of making a digital, searchable engine for historical American Newspapers (200, 2007). While previous studies have leveraged the Chronicling America collection, their focus has predominantly been on visual content extraction (Lee et al., 2020). In contrast, we focus on utilizing the already digitized text, and we employ this resource for testing cutting-edge information processing technologies on this novel and highly challenging type of text data.
@@
We adopt an automatic approach to construct the dataset due to several reasons. First, the historical document collection is huge, and processing such a sheer collection of data manually is impossible; because of this, we use GPT 3.5 (Brown et al., 2020) for correcting the digitized text available on Chronicling America. Second, we employ a generative model to generate the questions automatically because manual generation would be very costly. Third, recently developed generative models have shown great performance for automatic question generation (Du et al., 2017; Pan et al., 2019) at the same time. The final dataset that we release comprises 485k question-answer pairs spanning 120 years (1800-1920) [^2] The dataset is freely available at: https://github.com/DataScienceUIBK/ChroniclingAmericaQA.git .
@@
  • We quantify the performance degradation of various models for question-answering tasks due to Noisy OCR text.
@@ 2.2. QA Benchmarks
In recent years, numerous large-scale QA datasets have been introduced, contributing significantly to the field of QA. Notable examples include SQuAD 1.1 (Rajpurkar et al., 2016), a landmark dataset for reading comprehension derived from Wikipedia articles. It was followed by SQuAD 2.0 (Rajpurkar et al., 2018), which expanded the challenge by incorporating unanswerable questions. Beyond Wikipedia, datasets like NaturalQA (Kwiatkowski et al., 2019) and MS MARCO (Bajaj et al., 2016) have utilized queries issued to Google and Bing search engines as questions, top-ranking Wikipedia pages or other web documents as context. NarrativeQA (Kočiský et al., 2018) broadened the scope by using books and movie script summaries as the basis for its question-answer pairs.
@@
Thus, this work aims to create a QA dataset from the Historical Newspapers collection to foster the research of QA on Historical documents. Our dataset, ChroniclingAmericaQA, is designed to handle the unique challenges posed by historical newspapers, such as noisy OCR data, archaic language as well as the overall context of the distant time that differs from the present one [^3] Consider the term Zeitgeist in German language used to emphasize the spirit or essence of the particular time period or epoch.. We believe that this dataset will be a valuable resource for researchers in the field of QA and will contribute to the advancement of QA for historical documents and the NLP field in general.
@@ 3.1. Data Collection Module
We chose the newspaper pages from the collection of the American Historical Newspaper dataset Chronicling America [^4] https://chroniclingamerica.loc.gov/about/ available in the public domain as our source for question generation. Chronicling America is a project developed by the Library of Congress and the National Endowment for the Humanities (NEH) under the National Digital Newspaper Program (NDNP). It features over 21 million pages of historical American newspapers published from 1756 to 1963, accessible through an online search portal and public API. The portal offers various search strategies to access the content available on Chronicling America. The search can be performed in three ways: by identifying newspaper pages published in a specific state and year, conducting more advanced searches for newspaper pages containing particular words or phrases, and locating newspaper pages related to specific ethnic groups available in different languages. We opted for the first search strategy, i.e., statewide search by year, as our goal was to create a corpus of historical newspapers that covers diverse information ranging from every significant past event and ethnicity to some minor regional events throughout the United States. Therefore, we decided to search for the newspaper pages published across the 53 states in the United States. Figure 3 shows the time intervals for which newspapers are available for each state of the United States in the Chronicling America Collection. Considering the availability of data from different states, we opted for the period from 1800 to 1920, which spans 120 years.
@@ 3.2. Data Preparation Module
To address this issue, we post-process the raw OCR text to enhance its quality, thereby making it more suitable for question generation. Post-processing OCR text is challenging, as demonstrated by several studies (Nguyen et al., 2021; Jatowt et al., 2019). Traditional methods for post-OCR error correction, such as statistical language modeling techniques (Afli et al., 2015), lexical approaches (Strohmaier et al., 2003), merging OCR outputs (Lund and Ringger, 2009), are inadequate for handling historical text due to the unique complexities associated with such data.
@@
Before giving the raw OCR text to the GPT model for correction, we need to split the content of newspaper pages into paragraphs. Since the content on a single newspaper page is diverse and long, ranging from advertisements and articles to cartoons, processing such long text will be difficult for question generation as well as for correcting errors in the OCRed text. Therefore, we split the content a newspaper page into multiple paragraphs. We set the length of a paragraph to 250 words due to the context length limitation imposed by GPT 3.5 Turbo model [^6] The maximum output token length of gpt-3.5-turbo-1106 is 4096..
@@
In this manner, we tasked GPT 3.5 with correcting 205,638 paragraphs. Out of these, GPT was able to correct 163,021. We eliminated the uncorrected paragraphs (42,599) from our corpus and utilized the corrected paragraphs for question generation. Among the corrected paragraphs, 433 contained phrases such as ”Most of the mistakes in the given text have been corrected. Here is the corrected text:” and ”Errors in the given text have been corrected. Here is the corrected text:”. We removed such 96 different phrases, cleaned the paragraphs, and prepared them for question generation. Figure 4 illustrates the number of corrected paragraphs per decade.
@@ 3.3. Question Generation Module
Question & answer cleaning: Given that we employ the T5-base model, a generative model for question generation, the produced questions may exhibit various issues, such as duplication, explicit answer disclosure within the question, unresolved pronouns, and more. To address these issues, we implemented a multi-step filtering process to refine the dataset. The process includes:
@@
    Syntactic Filtering: In this initial step, we perform five key actions to refine the generated question-answer pairs. We start by eliminating questions that do not conclude with a question mark, followed by removing questions that reveal the answer within the question itself (answer leakage) and eliminating duplicates. We also filter out questions that are excessively long or short, as well as those with too many or too few number of named entities. Additionally, questions with unclear pronouns are removed. This step resulted in removing 1,221,533 questions from our initial pool of generated questions.
@@
    Type Matching: The third and final filtering step, involved identifying and matching the types of questions and answers. For this purpose, we finetined RoBERTa (Liu et al., 2019)- a large language model using the TREC Question classifier dataset (Li and Roth, 2002) to develop a question type classifier [^8] The accuracy of the question type classifier is 92.8%.. The question classifier model classifies the question type into four major types and 50 minor types. Similarly, for answer type classification, we utilized spaCy’s named entity recognizer. After identifying the question and answer type, we then matched the question-answer pairs. Question and answer pairs whose type was not the same were eliminated. In this step, we eliminated 1,205,482 question-answer pairs.
@@ 4.1. Data Statistics
After applying all the filtering steps, we arrived at our cleaned question-answer dataset consisting of 485K question-answer pairs derived from historical American newspapers. In Figure 5, we show the distribution of newspaper pages in our dataset across various states of America. We randomly split the data into training, development, and test sets to facilitate a comprehensive training and evaluation of the models. The training set consists of 439,302 question-answer pairs, whereas the development and test sets comprise 24,111 and 24,084, respectively. When randomly selecting the questions for development and test sets, we tried to balance the question-answer pairs across all the decades. We describe more details about the dataset statistics in Table 3. Our entire train set is quite large, which makes it computationally expensive to fine-tune various pre-trained models such as BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), T5 (Raffel et al., 2019); therefore, we randomly select a subset of question-answer (111,517) pairs from the train set and use this subset of questions for training various models whose results are shown in Section 4.2.
@@ 4.2. Model Performance
Table 5 summarizes the performance of different transformers on the ChroniclingAmericaQA dataset using standard QA metrics: exact match (EM) and F1 score. We measure how well different models can answer questions from both Corrected and Raw OCR paragraphs. From the results, it is evident that models used without any fine-tuning performed the worst, as these models lack specific knowledge about the question-answering task. Conversely, models fine-tuned on the SQuAD 1.1 dataset show significant improvements, highlighting the benefits of fine-tuning on task-specific datasets. Notably, the highest performance levels are observed from the models trained on the ChroniclingAmericaQA dataset itself, depicting the importance of fine-tuning on specific datasets. However, models trained on both SQuAD and ChroniclingAmericaQA give even better performance, suggesting that combining domain-specific knowledge such as historical content and general task specific knowledge can even enhance the performance of models.
We also compare the results of transformer models on both Raw OCR paragraphs and Corrected OCR paragraphs in Table 5. From the results, we can analyze that model performance significantly drops when the Raw OCR paragraph is given as context, suggesting the importance of cleaning the digitized text. We also show the performance of fine-tuned models over time in Figure 8. It can be concluded from the figure that the performance improvement due to finetunning on our dataset remains relatively stable over time.
In summary, we can conclude that results shown in table 5 ensure fine-tuning models on relevant datasets significantly improve the performance. Additionally, we can also realize that correcting the OCRed text is crucial for getting better model performance.
@@ 4.3. LLM Performance
Similarly, in addition to transformer models, we evaluate our dataset with Large Language Models (LLMs), including the LLaMA2 (tou, 2023) family, Mixtral 8x7B (Jiang et al., 2024) and Mistral 7B (Jiang et al., 2023). Evaluating the performance of LLMs (Guo et al., 2023) within the context of question answering poses unique challenges. Traditionally, QA systems are assessed using Exact Match (EM) and F1 scores. However, these metrics are not fully appropriate for generative models such as LLMs, which often produce verbose responses. This verbosity can result in an EM value of zero due to the presence of extra tokens in the answer. Similarly, the F1 score, which evaluates both recall and precision, may not provide an accurate measure of performance under these conditions. To overcome the challenge of evaluating LLMs, we introduce model-agnostic metrics like Token Recall and Answer String Containment [^15] https://huggingface.co/spaces/evaluate-metric/squad, which are more suitable for assessing verbose responses. Token Recall evaluates how effectively the model’s response covers the information found in the ground truth, while Answer String Containment assesses the degree to which the ground truth is included within the model’s response. By using these metrics, we aim to offer a more equitable evaluation of an LLM’s ability to generate responses to questions.
Table 6 presents the performance of different LLMs on the ChroniclingAmericaQA Dataset using both raw OCR and corrected paragraphs. When evaluating the results of different LLMs, we can note that LLaMA2 70B gives significantly high EM and F1 scores compared to its counterparts and Mistral models. This also suggests that models with large sizes have a greater ability to understand the nuances of historical text, leading to more accurate and precise answers. Meanwhile, the Mixtral 8x7b model performs highly on Recall and Contains, depicting its proficiency in extracting relevant information from the context.
Furthermore, When examining the results of different LLMs considering the raw OCR paragraph and corrected paragraph, we can note that performance across all models decreased drastically when utilizing the raw OCR paragraph as context. Models such as LLaMA2 7B and Mixtral Bx7B gave zero results on EM metric, whereas LLaMA2 70B still showed some resistance on raw OCR paragraph, however, with a significant drop of 93%. Furthermore, When analyzing the recall and precision metrics results, one can note that models showed better performance on recall, indicating their ability to capture relevant information from the context.
Summarizing the results of Table 6, we can conclude that the performance of LLaMA2 70B is high compared to other models for handling both raw OCR text and corrected text. On the other hand, Mixtral 8x7B gives high results for recall and contains metrics on corrected OCR paragraphs. However, LLaMa2 70B gives a high recall and contains value on the raw OCR paragraph.

### v2

@@ Abstract.
Question answering (QA) and Machine Reading Comprehension (MRC) tasks have significantly advanced in recent years due to the rapid development of deep learning techniques and, more recently, large language models. At the same time, many benchmark datasets have become available for QA and MRC tasks. However, most existing large-scale benchmark datasets have been created predominantly using synchronous document collections like Wikipedia or the Web. Archival document collections, such as historical newspapers, contain valuable information from the past that is still not widely used to train large language models. To further contribute to advancing QA and MRC tasks and to overcome the limitation of previous datasets, we introduce ChroniclingAmericaQA, a large-scale temporal QA dataset with 487K question-answer pairs created based on the historical newspaper collection Chronicling America. Our dataset is constructed from a subset of the Chronicling America newspaper collection spanning 120 years. One of the significant challenges for utilizing digitized historical newspaper collections is the low quality of OCR text. Therefore, to enable realistic testing of QA models, our dataset can be used in three different ways: answering questions from raw and noisy content, answering questions from cleaner, corrected version of the content, as well as answering questions from scanned images of newspaper pages. This and the fact that ChroniclingAmericaQA spans the longest time period among available QA datasets make it quite a unique and useful resource.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3657891isbn: 979-8-4007-0431-4/24/07ccs: Information systems Question answeringccs: Information systems Content analysis and feature selection
@@ 1. Introduction
However, leveraging these documents for QA is complex and poses a few challenges. The language used in heritage documents often differs significantly from the contemporary language in terms of vocabulary, syntax, and the overall context of the distant times (Tahmasebi et al., 2021). These raises doubts if QA models trained on current text data can indeed perform well on documents from distant past. Additionally, the documents are often available only in scanned form, and the quality of the text obtained through Optical Character Recognition (OCR) can greatly vary, especially over long time periods, adding another layer of complexity to the task. Finally, heritage news articles have a complex layout, often containing different font sizes and type throughout their content, which makes information extraction a complex task. Figure 1 shows the example of a scanned newspaper page and its OCRed text.
@@
Figure 1. An example of the scanned newspaper page from Chronicling America Collection. a) depicts the entire newspaper page published in Evening Star on 1803-02-07 in the District of Columbia, b) depicts the zoomed-in paragraph of the newspaper page shown in (a), c) shows the original OCR text of the zoomed-in paragraph that is available in the Chronicling America, and d) displays the OCR text corrected by GPT 3.5 Turbo.
@@
Despite the inherent challenges of using historical newspapers, they present an attractive and unique opportunity for research in the field of QA. The wealth of information in such documents makes them valuable resources that are not readily available in typical evaluation benchmarks. To address the challenges and utilize the potential of historical document collections, we introduce the ChroniclingAmericaQA, a large-scale dataset based on Historical American Newspaper Pages. As an underlying historical document collection, we use Chronicling America [^1] https://chroniclingamerica.loc.gov/about/, a digitized collection of America’s Historical newspaper pages from 1756-1963. Chronicling America is a project that was created with the aim of making a digital, searchable engine for historical American Newspapers (Culpepper, 2007). While previous studies have leveraged the Chronicling America collection, their focus has predominantly been on visual content extraction (Lee et al., 2020). In contrast, we focus on utilizing the already digitized text, and we employ this resource for testing cutting-edge information processing technologies to answer natural language questions from this novel and highly challenging type of text data.
@@
We adopt an automatic approach to construct the dataset due to several reasons. First, the historical document collection is huge, and processing such a sheer collection of data manually is impossible. Because of this, we use GPT 3.5 (Brown et al., 2020) for correcting the digitized text available on Chronicling America. Second, we employ a generative model to generate the questions automatically because manual generation would be very costly. Third, recently developed generative models have shown great performance for automatic question generation (Du et al., 2017; Pan et al., 2019). The final dataset that we release comprises 487K question-answer pairs spanning 120 years (1800-1920) [^2] The dataset is freely available at: https://github.com/DataScienceUIBK/ChroniclingAmericaQA .
@@
  • We quantify the performance degradation of various models for question-answering tasks caused by noisy OCR text.
@@ 2.2. QA Benchmarks
In recent years, numerous large-scale QA datasets have been introduced, contributing significantly to the field of QA. Notable examples include SQuAD 1.1 (Rajpurkar et al., 2016), a landmark dataset for reading comprehension derived from Wikipedia articles. It was followed by SQuAD 2.0 (Rajpurkar et al., 2018), which expanded the challenge by incorporating unanswerable questions. Beyond Wikipedia, datasets like NaturalQA (Kwiatkowski et al., 2019) and MS MARCO (Bajaj et al., 2016) have utilized queries issued to Google and Bing search engines as questions, top-ranking Wikipedia pages or other web documents as context. NarrativeQA (Kočiský et al., 2018) broadened the scope by using books and movie script summaries as the basis for its question-answer pairs, and HotpotQA (Yang et al., 2018) further extended the task to multi-hop question answering (Mavi et al., 2022).
@@
Thus, this work aims to create a QA dataset from the historical newspapers collection to foster the research of QA on historical documents. Our dataset, ChroniclingAmericaQA, is designed to handle the unique challenges posed by historical newspapers, such as noisy OCR data, archaic language as well as the overall context of the distant time that differs from the present one [^3] For example, consider the term Zeitgeist in German used to emphasize the spirit or essence of the particular time period or epoch.. We believe that this dataset will be a valuable resource for researchers in the field of QA and will contribute to the advancement of QA for historical documents and the NLP field in general.
@@ 3.1. Data Collection Module
We chose the newspaper pages from the collection of the American historical newspaper dataset, Chronicling America, [^4] https://chroniclingamerica.loc.gov/about/ available in the public domain as our source for question generation. Chronicling America is a project developed by the Library of Congress and the National Endowment for the Humanities (NEH) under the National Digital Newspaper Program (NDNP). It features over 21 million pages of historical American newspapers published from 1756 to 1963, accessible through an online search portal and public API. The portal offers various search strategies to access the content available on Chronicling America. The search can be performed in three ways: by identifying newspaper pages published in a specific state and year, conducting more advanced searches for newspaper pages containing particular words or phrases, and locating newspaper pages related to specific ethnic groups available in different languages. We opted for the first search strategy, i.e., statewide search by year, as our goal was to create a corpus of historical newspapers that covers diverse information ranging from every significant past event and ethnicity to some minor regional events throughout the United States. Therefore, we decided to search for the newspaper pages published across the 53 states in the United States. Figure 3 shows the time intervals for which newspapers are available for each state of the United States in the Chronicling America Collection. Considering the availability of data from different states, we opted for the period from 1800 to 1920, which spans 120 years.
@@ 3.2. Data Preparation Module
To address this issue, we post-process the raw OCR text to enhance its quality, thereby making it more suitable for question generation. Post-processing OCR text is challenging, as demonstrated by several studies (Nguyen et al., 2021; Jatowt et al., 2019). Traditional methods for post-OCR error correction, such as statistical language modeling techniques (Afli et al., 2016), lexical approaches (Strohmaier et al., 2003), merging OCR outputs (Lund and Ringger, 2009), often fall short when dealing with historical texts due to the unique complexities associated with such data.
@@
Before giving the raw OCR text to the GPT model for correction, we need to split the content of newspaper pages into paragraphs. Since the content on a single newspaper page is typically diverse and long, ranging from advertisements and articles to cartoons, processing such long text will be difficult for question generation as well as for correcting errors in the OCRed text. Therefore, we split the content a newspaper page into multiple paragraphs. We set the length of a paragraph to 250 words due to the context length limitation imposed by GPT 3.5 Turbo model [^6] The maximum output token length of gpt-3.5-turbo-1106 is 4096..
@@
In this manner, we tasked GPT 3.5 with correcting 205,638 paragraphs. Out of these, GPT was able to correct 163,021. We eliminated the uncorrected paragraphs (42,599) from our corpus and utilized the corrected paragraphs for question generation. Among the corrected paragraphs, 433 contained phrases such as ”Most of the mistakes in the given text have been corrected. Here is the corrected text:” and ”Errors in the given text have been corrected. Here is the corrected text:”. We removed 96 patterns of such different phrases, cleaned the paragraphs, and prepared them for question generation. Figure 4 illustrates the number of corrected paragraphs per decade.
@@ 3.3. Question Generation Module
Question & answer cleaning: Given that we employ the T5-base model, a generative model for question creation, the produced questions may have various issues, such as duplication, explicit answer disclosure within the question, unresolved pronouns, and more. To address these issues, we implemented a multi-step filtering process to refine the dataset. The process includes:
@@
    Syntactic Filtering: In this initial step, we perform five key actions to refine the generated question-answer pairs. We start by eliminating questions that do not conclude with a question mark, followed by removing questions that reveal the answer within the question itself (answer leakage) and eliminating duplicates. We also filter out questions that are excessively long or short, as well as those with too many or too few number of named entities. Additionally, questions with unclear pronouns, for example, ”Who was General of the United States at the time of your letter to me?” are removed. This step resulted in removing 1,221,533 questions from our initial pool of generated questions.
@@
    Type Matching: The third and final filtering step, involved identifying and matching the types of questions and answers. For this purpose, we finetined RoBERTa (Liu et al., 2019)- a large language model using the TREC Question classification dataset (Li and Roth, 2002) to develop a question type classifier [^8] The accuracy of the question type classifier is 92.8%.. The question classifier model classifies the question type into four major types and 50 minor types. Similarly, for answer type classification, we utilized spaCy’s named entity recognizer. After identifying the question and answer type, we then matched the question-answer pairs. Question and answer pairs whose type was not the same were eliminated. In this step, we eliminated 1,205,482 question-answer pairs.
@@ 4.1. Data Statistics
After applying all the filtering steps, we arrived at our cleaned question-answer dataset consisting of 487K question-answer pairs derived from historical American newspapers. In Figure 5, we show the distribution of newspaper pages in our dataset across the states of America. We randomly split the data into training, development, and test sets to facilitate a comprehensive training and evaluation of the models. The training set consists of 439,302 question-answer pairs, whereas the development and test sets comprise 24,111 and 24,084, respectively. When randomly selecting the questions for development and test sets, we tried to balance the question-answer pairs across all the decades. We describe more details about the dataset statistics in Table 3. Our entire train set is quite large, which makes it computationally expensive to fine-tune various pre-trained models such as BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), T5 (Raffel et al., 2019); therefore, we randomly select a subset of question-answer (111,517) pairs from the train set and use this subset of questions for training various models whose results are shown in Section 4.2.
@@ 4.2. Model Performance
Table 5 summarizes the performance of different transformers on the ChroniclingAmericaQA dataset using standard QA metrics: Exact Match (EM) and F1 score. We measure how well different models can answer questions from both Corrected and Raw OCR paragraphs. From the results, it is evident that models used without any fine-tuning performed the worst, as these models lack specific knowledge about the question-answering task. Conversely, models fine-tuned on the SQuAD 1.1 dataset show significant improvements, highlighting the benefits of fine-tuning on task-specific datasets. Notably, the highest performance levels are observed from the models trained on the ChroniclingAmericaQA dataset itself, depicting the importance of fine-tuning on specific datasets. However, models trained on both SQuAD and ChroniclingAmericaQA give even better performance, suggesting that combining domain-specific knowledge such as historical content and general task specific knowledge can enhance the performance of models.
We also compare the results of transformer models on both Raw OCR paragraphs and Corrected OCR paragraphs in Table 5. From the results, we can see that model performance significantly drops when the Raw OCR paragraph is given as context, suggesting the importance of cleaning the digitized text. We also show the performance of fine-tuned models over time in Figure 8. It can be concluded from the figure that the performance improvement due to finetunning on our dataset remains relatively stable over time.
In summary, we can conclude that results shown in Table 5 imply that fine-tuning models on relevant datasets significantly improves the performance. Additionally, we can also conclude that correcting the OCRed text is crucial for getting better model performance.
@@ 4.3. LLM Performance
Similarly, in addition to transformer models, we evaluate our dataset with Large Language Models (LLMs), including the LLaMA2 (Touvron et al., 2023) family, Mixtral 8x7B (Jiang et al., 2024) and Mistral 7B (Jiang et al., 2023). Evaluating the performance of LLMs (Guo et al., 2023) within the context of question answering poses unique challenges. Traditionally, QA systems are assessed using Exact Match (EM) and F1 scores. However, these metrics are not fully appropriate for generative models such as LLMs, which often produce verbose responses. This verbosity can result in an EM value of zero due to the presence of extra tokens in the answer. Similarly, the F1 score, which combines both recall and precision, may not provide an accurate measure of performance under these conditions. To overcome the challenge of evaluating LLMs, we introduce model-agnostic metrics like Token Recall and Answer String Containment [^15] https://huggingface.co/spaces/evaluate-metric/squad, which are more suitable for assessing verbose responses. Token Recall evaluates how effectively the model’s response covers the information found in the ground truth, while Answer String Containment assesses the degree to which the ground truth is included within the model’s response. By using these metrics, we aim to offer a more equitable evaluation of an LLM’s ability to generate responses to questions.
Table 6 presents the performance of different LLMs on the ChroniclingAmericaQA Dataset using both raw OCR and corrected paragraphs. When evaluating the results of different LLMs, we can note that LLaMA2 70B gives significantly higher EM and F1 scores compared to its counterparts and Mistral models. This also suggests that models with large sizes have a greater ability to understand the nuances of historical text, leading to more accurate and precise answers. Meanwhile, the Mixtral 8x7b model performs well on recall and contains metrics, depicting its proficiency in extracting relevant information from the context.
Furthermore, when examining the results of different LLMs considering the raw OCR paragraphs and corrected paragraphs, we notice that performance across all models decreased drastically when utilizing the raw OCR paragraphs as context. Models such as LLaMA2 7B and Mixtral Bx7B gave zero results on EM metric, whereas LLaMA2 70B still showed some resistance on raw OCR paragraphs, however, with a significant drop of 93%. Moreover, when analyzing the recall and precision metrics results, one can notice that models showed better performance on recall, indicating their ability to capture relevant information from the context.
Summarizing the results of Table 6, we can conclude that the performance of LLaMA2 70B is better compared to other models when handling both raw OCR text and corrected text. On the other hand, Mixtral 8x7B gives good results for recall and contains metrics on corrected OCR paragraphs. However, LLaMa2 70B has a high recall and contains values on the raw OCR paragraphs.